OpenStudy (anonymous):

find the area of the triangle with vertices at (1,0), (0,-1) and (2,-2)

OpenStudy (anonymous):

first find length of each side then apply heron's formulae

OpenStudy (anonymous):

the sides are \[\sqrt{2}, \sqrt{5}, \sqrt{5}?\] right? but it's quite complicated to use the heron's formulae for this. is there another method?

OpenStudy (anonymous):

there is formula det ([x1 x2 1] [y1 y2 1 ] [z1 z2 1 ] ) just plug your coordinates into this formula and get your ans

OpenStudy (anonymous):

1/2 times above formula

OpenStudy (anonymous):

is there another way other than using a matrix?

OpenStudy (anonymous):

3/2~~1.5

OpenStudy (anonymous):

ya u have a method other than matrices

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

This is an isosceles triangle...

OpenStudy (anonymous):

 Here is a different approach: - the triangle fits in the box drawn by the point (2, -2), so the area is less than 4 (= 2 * 2) - Inside this (2,-2) box, there are 4 triangles after drawing the triangle we want to know the area of. - the area of the other 3 triangles is easy to calculate because they are all right-triangles A = base * height / 2 A_triangle1 = 1*1/2 = 1/2 A_triangle2 = 1*2/2 = 1 A_triangle3 = 2*1/2 = 1 A_triangle4 = this is the one we want to figure out 4 = A_triangle1 + A_triangle2 + A_triangle3 + A_triangle4 4 = 2.5 + A_triangle4 A_triangle4 = 1.5